OpenStudy (anonymous):

sert y=0 and solve 4x- (12-4x)^1/2

OpenStudy (anonymous):

y = 4x - (12-4x)^1/2 if y = 0 \[0= 4x - \sqrt{12-4x}\] \[\sqrt{12-4x} = 4x\] square both sides \[12-4x = 16 x ^{2}\] \[12 = 16x ^{2} + 4x\] then divide by 16 \[\frac{ 3 }{ 4 } = x ^{2} +\frac{ 1 }{ 4 }x\] completing the squares \[\frac{ 3 }{ 4 } + \frac{ 1 }{ 64 } = (x + \frac{ 1 }{ 8 })^{2}\] \[\sqrt{\frac{ 49 }{ 64 }} = x + \frac{ 1 }{ 8 }\] \[-\frac{ 1 }{ 8 } \pm \frac{ 7 }{ 8 } = x\] x can be x= -1 x = 3/4

OpenStudy (anonymous):

correct me if im wrong to something sorry i have no papers here and im on a shop

OpenStudy (anonymous):

yup x=-1 or x=3/4. But x=3/4 is the only answer where 4(3/4) - (12-4(3/4))^1/2 = 0 which satisfies with your y = 0 But when x=-1 4(-1) - (12-4(-1))^1/2 = -8 where it fails -8 is not equal to 0.. So x=3/4 is the only answer